OpenStudy (kainui):

Are there any other popular eigenfunctions other than e^x, sinx, cosx, and combinations of those? And before you say sinx isn't an eigenfunction, it is an eigenfunction of the second derivative operator.

OpenStudy (anonymous):

I don't believe so. I'm guessing a function that is constantly zero doesn't count...

OpenStudy (kainui):

Wow, I did some looking and found some really simple ways to make eigenfunctions. It's really all in the operator. For instance: \[x \frac{ d }{ dx }(x^n)=nx^n\] So you can pretty simply see that x^n is an eigenfunction of the x*d/dx operator. Kind of silly to see now that it's obvious. So for instance (lol) \[tan(ax )*\frac{ d }{ dx }[sin(ax)]=asin(ax)\] and now sin(ax) is an eigenfunction. Bleh. But is it useful for anything?

OpenStudy (anonymous):

isn't that sort of cheating?